BTGmoderatorLU wrote:In a high school debating team consisting of 2 freshmen, 2 sophomores, 2 juniors, and 2 seniors, two students are selected to represent the school at the state debating championship. The rules stipulate that the representatives must be from different grades, but otherwise the 2 representatives are to be chosen by lottery. What is the probability that the students selected will consist one freshman and one sophomore?
A. 1/16
B. 1/8
C. 1/7
D. 1/6
E. 1/4
Two of the four grades must be selected to represent the team at the championship.
Because each grade on the debate team offers the same number of students -- 2 students per grade -- the probability of picking each grade is the same.
Thus, the problem can be rephrased as follows:
What is the probability that the two grades selected are freshman and sophomore?
P(first grade selected is freshman or sophomore) = 2/4. (Of the 4 grades, 2 are either freshman or sophomore.)
P(second grade selected is freshman or sophomore) = 1/3. (Of the 3 grades not yet selected, 1 will be freshman or sophomore.)
To combine these probabilities, we multiply:
2/4 * 1/3 = 1/6.
The correct answer is
D.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.
As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.
For more information, please email me (Mitch Hunt) at
[email protected].
Student Review #1
Student Review #2
Student Review #3